OpenStudy (anonymous):

Aliyah left the white house and drove toward her cabin on the lake at an average speed of 45 mph. Gabriella left some time later driving in the opposite direction with an average speed of 70 mph. After aliyah had driven for five hours hey were 505 mi. apart. How long did Gabriella drive?

OpenStudy (anonymous):

Help me with this I don't know to do this s I can do this fro the rest.

OpenStudy (anonymous):

need the answer and how to do this

OpenStudy (koikkara):

hey, try these may be useful and save it....

OpenStudy (anonymous):

Thank you for that notes I will print it out

OpenStudy (anonymous):

But I kinda still don't know hwo to do this

OpenStudy (koikkara):

@Opcode ...help her..

OpenStudy (anonymous):

n how about this one Darryl traveled to the lake and back. It took one hour longer to go there than it did to come back. The average speed on the trip there was 36mph. The average speed on the way back was 45 mph. How many hours did the trip there take?

OpenStudy (anonymous):

r*t = d Let t be the time that Gabriella drove. Solve the following for t: 45*5 + 70 t = 505 t = 4 hours

OpenStudy (anonymous):

ok I get it thanks XD

OpenStudy (anonymous):

kropot stay

OpenStudy (anonymous):

A diesel train left the station and traveled west. One hour later a ccattle train left traveling at 22 mph in an effort to catch up to the diesel train. After traveling for 10 hours the cattle train finally caught up. Find the diesel train's average speed. Is it 220 as the speed. answer is 220?

OpenStudy (anonymous):

Also lok at the other one you helped me on

OpenStudy (kropot72):

The distance travelled by the cattle train = 22 * 10 = 220 miles The diesel train took 11 hours to travel the same distance (220 miles). Average speed of diesel train = 220/11 = ? mph

OpenStudy (anonymous):

where did you get 11 hoours from?

OpenStudy (anonymous):

oh you added 1 hour later when the cattle train started traveling ok XD